The fundamental equations in an economy are given as:

Consumption function C =200+0.8yd

Investment function. I=300

Tax. T=120

Government expenditure. G=200

Exports. X=100

Imports M=0.05y

Find the following.

1.The equilibrium level of income.

2.The net exports.



1
Expert's answer
2021-05-05T13:42:56-0400

The consumption function is given

"C=200+0.8yd"

"C=200+0.8(Y-T)"

"C=200+0.8(Y-120)"


The equilibrium is

"Y=C+I+G+X-M"

"=200+0.8(Y-120)+300+200(100-0.05Y)"

"200+0.8Y-96+600-0.05Y"

"Y-0.8Y+0.05Y=704"

"0.25Y=704"

"Y=\frac{704}{0.25}=2,816"


The equilibrium level of income is 2,816


Equilibrium in a four sector model, leakages is equal to injections

so, "C+I+G+X=C+S+T+M"

The consumption function is "C = 200 + 0.8Y"

"C=200+0.8(2,816-120)"

"C=200+2,156.8"

"C=2,356.8"


The saving function is

"S=Yd-C"

"=(Y-120)-2,356.8"

"=2,816-2476.8"

"S=339.2"


Thus,

"I + G + X = S + T + M"

"300 + 200 + 100 = 339.2 + 120 + 0.05Y"

"600 = 459.2 + 0.05 Y"

"600 = 459.2 + 0.05 (2,816)"

"600 = 600"


Imports      

"M = 0.05Y = 0.05 (2,816)"

"=140.8"

Net Exports:

"X \u2013 M = 100 \u2013 140.8"

"X - M = - 40.8"


there is deficit in the balance of trade.